anonymous202:

Dianna made several loaves of bread yesterday. Each loaf required 2 and two-thirds cups of flour. All together, she used 13 and one-third cups of flour. How many loaves did Dianna make? 3 loaves 4 loaves 5 loaves 7 loaves

anonymous202:

can someone please help I'm stuck on this

Oddmoyu123:

5 loaves you divide 13 1/3 by 2 2/3

Oddmoyu123:

trust jhonny over me tho

jhonyy9:

each loaf 2 2/3 cups of flour she used 13 1/3 cups 2 2/3 = 8/3 13 1/3 = 40/3 40/3 / 8/3 = 40/3 * 3/8 = 40/8 = 5

anonymous202:

thankyou

jhonyy9:

anytime i m glad to help